OpenStudy (londoncat):

I need to answer as an interval, help please, will medal Find all real numbers t such that (2/3 t )- 1 < t + 7 <= -2t + 15. I think that the answer is 8 > t < = 24/9, but i dont know how to submit it in the right format.

jimthompson5910 (jim_thompson5910):

Hint: Break the compound inequality \[\Large \frac{2}{3}t - 1 < t+7 \le -2t+15\] into these two inequalities \[\Large \frac{2}{3}t - 1 < t+7\] and \[\Large t+7 \le -2t+15\]

OpenStudy (xguardians):

I just realized it mean less than or equal to, my bad.

OpenStudy (londoncat):

How so you answer something in the form of an interval?

OpenStudy (freckles):

do you believe you have soled both inequalities correctly?

OpenStudy (freckles):

solved*

OpenStudy (freckles):

if so what did you get for both inequalities mentioned by jim?

OpenStudy (freckles):

assuming a is less than b: interval notation for a<x<b is (a,b) interval notation for a<=x<=b is [a,b] interval notation for a<=x<b is [a,b) interval notation for a<x<=b is (a,b] interval notation for x<a or x>b is (-inf,a) union (b,inf) interval notation for x<=a or x>b is (-inf,a] union (b,inf) interval notation for x<=a or x>=b is (-inf,a] union [b,inf) interval notation for x<a or x>=b is (-inf, a) union [b,inf) i hope this helps I can do some examples if you want actual numbers instead of a and b that represent some numbers where a is less than b

OpenStudy (londoncat):

im not entirely sure, im quite confused. After I separate the compound inequality what do I do?

OpenStudy (freckles):

solve both for t

OpenStudy (freckles):

let's play with the first mentioned one: \[\frac{2}{3}t-1<t+7\] if you don't like the fraction right now multiply on through by 3 giving \[2t-3<3t+21\]

OpenStudy (freckles):

your job is to isolate t

OpenStudy (freckles):

get your t-terms on one side and your non t-terms on the opposing side by addition or subtraction

OpenStudy (londoncat):

so 2/3t<t+8

OpenStudy (freckles):

ok that looks fine so far

OpenStudy (londoncat):

and for inequality 2, t<=-2t+8?

OpenStudy (freckles):

ok but you didn't finishing solving the first for t yet but yes so far RIGHT(as in correct) on inequality 2 which is still is unfinished also

OpenStudy (londoncat):

okay so to continue with 1, is it legal to say t<t+1/3t+8?

OpenStudy (freckles):

2/3 t < t+ 8 you want the t's on one side together subtract t on both sides

OpenStudy (freckles):

\[\frac{2}{3} t < t+8 \\ \text{ subtract} t \text{ on both sides } \\ \frac{2}{3}t\color{red}{-t}<t\color{red}{-t}+8 \\ \\ \frac{2}{3}t-t<0+8 \\ \frac{2}{3}t-t<8\]

OpenStudy (freckles):

\[\frac{2}{3}t-t<8 \\ \frac{2}{3}t-1 t<8 \\ (\frac{2}{3}-1)t<8 \]

OpenStudy (londoncat):

ohhhh okay...is that all for 1

OpenStudy (freckles):

no you have to solve for t

OpenStudy (freckles):

this means the only next to t can be 1 you must solve for t aka 1t

OpenStudy (londoncat):

oh duh whoops

OpenStudy (freckles):

you have (2/3-1) next to t which is not 1

OpenStudy (londoncat):

so 1>t<8

OpenStudy (freckles):

how did you that?

OpenStudy (freckles):

\[\frac{2}{3} t < t+8 \\ \text{ subtract} t \text{ on both sides } \\ \frac{2}{3}t\color{red}{-t}<t\color{red}{-t}+8 \\ \\ \frac{2}{3}t-t<0+8 \\ \] \[\frac{2}{3}t-t<8 \\ \frac{2}{3}t-1 t<8 \\ (\frac{2}{3}-1)t<8 \] do you know how to do 2/3-1 ?

OpenStudy (londoncat):

-1/3?

OpenStudy (freckles):

right \[-\frac{1}{3} t<8\]

OpenStudy (freckles):

lastly for the first inequality you need to multiply both sides by -3

OpenStudy (freckles):

what inequality do you have after doing that?

OpenStudy (londoncat):

so t=24 for number 1

OpenStudy (londoncat):

-24

OpenStudy (londoncat):

wait

OpenStudy (freckles):

not exactly you dropped your inequality part and replaced it with an equality sign ok the -24 part is right what sign should you have between t and -24 though ?

OpenStudy (londoncat):

t<-24

OpenStudy (londoncat):

there

OpenStudy (freckles):

so if 2<3 then you are saying -2<-3 (I multiplied both sides by a negative)?

OpenStudy (freckles):

what I'm saying is if you multiply or divided both sides by a negative number on an inequality then the sign flips

OpenStudy (freckles):

\[\frac{-1}{3} t<8 \\ t>8(-3) \\ t>-24\]

OpenStudy (londoncat):

oh right, because we are multiplying by neg t would actually be bigger just like -2 is bigger than -3, so what i actually mean is t>-24

OpenStudy (freckles):

yah!

OpenStudy (londoncat):

now that we have this t we solve inequality 2 the same right?

OpenStudy (freckles):

\[\Large t+7 \le -2t+15 \\ \text{ here you already subtracted 7 on both sides } \\ t \le -2t +8\]

OpenStudy (freckles):

isolate t again

OpenStudy (freckles):

you figure out a way to some how get all your t-terms on the same side

OpenStudy (londoncat):

so 3t<=8?

OpenStudy (freckles):

beautiful so far you added 2t on both sides great deal

OpenStudy (londoncat):

so t<=8/3

OpenStudy (freckles):

right

OpenStudy (freckles):

so we want all numbers t such that \[t>-24 \text{ and } t \le \frac{8}{3}\]

OpenStudy (freckles):

can you write this as a compound inequality?

OpenStudy (londoncat):

[-25,8/3)

OpenStudy (londoncat):

oh

OpenStudy (freckles):

I bet you mean (-24,8/3] we want to include 8/3 but not include -24 the brackets means to include (think equal here) the parenthesis means to exclude ( think no equal there)

OpenStudy (londoncat):

-24<t<=8/3

OpenStudy (londoncat):

oh yeah:) thanks

OpenStudy (londoncat):

so is that all?

OpenStudy (freckles):

yep the answer in inequality notation is -24<t<=8/3 while in interval notation (-24,8/3]

OpenStudy (londoncat):

Thank you sososososososososo much!!!!!!!!!!!!!

OpenStudy (freckles):

np